OpenStudy (lolipop_drum312):

Can someone Please HELP me with 10 Questions i don'ti don't want to fail my test and i don't have much time left. All the Q's i will post will probably not have an A-D thing but i will try to add it.

OpenStudy (lolipop_drum312):

Which ordered pairs lie on the graph of the exponential equation ​xf(x)=128(0.5)x ​ ? Select each correct answer. ​A. (0,1) ​ B. (1,64) ​C. (3,16) ​​ D. (8,0.5)

OpenStudy (lolipop_drum312):

What function represents the graph y=4^x−1y? A. q B. r C. s D. t https://static.k12.com/nextgen_media/assets/8078551-NG_AL1_SemA_06_UT_40.png

OpenStudy (lolipop_drum312):

The function f(x)=(25^)xis shown on the coordinate plane. Select from the drop-down menus to correctly describe the end behavior of f(x)f(x) . As x decreases without bound, the graph of f(x)f(x) . As x increases without bound, the graph of f(x)f(x) .

OpenStudy (snoopdogg):

ill help but what do you think number 1 is

OpenStudy (mathmale):

I hope these questions are not straight from a test. In that case, we cannot help you.

OpenStudy (mathmale):

I don't believe that your "xf(x)=128(0.5)x" in Question 1 is correct. Was there really an "x" in front of "f(x)"?

OpenStudy (mathmale):

Also, the right side should be 128(0.5)^x. The " ^ " sign indicates "exponentiation" and is required.

OpenStudy (snoopdogg):

wow you can help him but not me soo unfair